Write the equation of a circle with center (4, 2) and radius r = 7.
(x – 2)^2 + (y – 4)^2 = 49
(x – 4)^2 + (y – 2)^2 = 49
(x + 4)^2 + (y + 2)^2 = 49
(x + 2)^2 + (y + 4)^2 = 49

Answer by ikleyn(52781)

.
Write the equation of a circle with center (4, 2) and radius r = 7.

(x – 2)^2 + (y – 4)^2 = 49

(x – 4)^2 + (y – 2)^2 = 49     <<<---===  This one.

(x + 4)^2 + (y + 2)^2 = 49

(x + 2)^2 + (y + 4)^2 = 49


It is the same as to ask "what is 2*2 ?".

Learn from your textbook.